# The Curious Case of Arbitrariness in Machine Learning

### v1

@@ 2.2. Impact of Rashomon Effect
The most straightforward example of multiplicity is when individuals encounter conflicting predictions from models in the Rashomon set, known as predictive multiplicity (Marx et al., 2020). Such conflicts create brittleness and arbitrariness in decision-making, undermining the reliability of these models and hampering effective planning (D’Amour et al., 2022; Marx et al., 2020; Black et al., 2022; Cooper et al., 2022; Watson-Daniels et al., 2024; Milani Fard et al., 2016). While predictive multiplicity can be harmful in critical domains, such as medical or legal decisions, it is, however, not inherently bad. For instance, purposefully controlled arbitrariness (called ‘randomness’ to distinguish from uncontrolled arbitrariness) in predictions can help address the concerns of outcome homogenization (Jain et al., 2024b, a; Creel and Hellman, 2022; Barocas et al., 2023; Perry and Zarsky, 2015).
@@
Unsurprisingly, predictive multiplicity has received significant attention in the literature ($35\%$ of papers in our systematic review; see Figure 1). Yet, this is only one aspect of the Rashomon effect. Consider, for instance, the rapidly growing field of inconsistency in explanations provided by models within the Rashomon set ($37.5\%$ of papers in our systematic review; see Figure 1). Studies have shown that models in the Rashomon set often produce conflicting variable importance and feature attribution scores (Black et al., 2021; Laberge et al., 2023; Hancox-Li, 2020; Baniecki et al., 2024; Li et al., 2024; Gunasekaran et al., 2024; Müller et al., 2023; Poiret et al., 2023; Okazaki et al., 2024; Tan, 2024; Kuzilek and Çavuş, 2024; Kumar et al., 2023), which can undermine trust, for example, confusing clinicians that rely on explanations during AI-assisted diagnostics. Similarly, counterfactuals generated by one model in the Rashomon set often fail to transfer to others (Pawelczyk et al., 2020; Leofante et al., 2023; Jiang et al., 2024a; Hasan and Talbert, 2022). This poses significant challenges for algorithmic recourse when models are regularly updated, as recourse provided by one model may become invalid when replaced by another, questioning their legitimacy and undermining user trust (Rawal et al., 2021; Leofante et al., 2023).
@@
More broadly, the effects of the Rashomon effect—multiplicity—can be measured as any form of behavioural difference between models in the context of real-world impact. These effects are multifaceted, and while predictive and explanation multiplicities are the most recognized, narrowing our focus underestimates the broader risks associated with other underrepresented forms of multiplicity (only $\sim 24\%$ of papers in our systematic review cover other forms of multiplicity; see Figure 1). In the next section, we will formally define both Rashomon sets and multiplicity, expanding upon existing definitions in the literature and drawing inspiration from the discussion presented here.
3. Definitions and Metrics
@@
We now define Rashomon sets and multiplicity, building on existing literature while expanding the scope to encompass a wider range of works. Additionally, we review multiplicity metrics and the literature on evaluating multiplicity.
3.1. Formalizing Rashomon Sets and Multiplicity
@@
The concept of multiplicity is deeply rooted in the Rashomon effect, and the models illustrating this Rashomon effect are together known as a Rashomon set, a set of competing models, a set of good models, $\epsilon$-Rashomon set, $\epsilon$-Level set, etc. We’ll stick with the term Rashomon set for consistency in our formalization. In essence, the Rashomon set represents a set of models that are practically indistinguishable, underscoring the arbitrariness in choosing one model over another. Thus, we need to begin by defining these models and explaining what it means for them to be indistinguishable.
@@
Existing work in multiplicity tends to adopt a narrow view of these models. For instance, much of the research that formalizes the multiplicity problem restricts model choices to a specific hypothesis class $\mathcal{H}$ and/or limits them to training on a fixed, pre-processed dataset $\mathcal{D}$ (Marx et al., 2020; Castillo et al., 2008; Teney et al., 2022; Watson-Daniels et al., 2023b, a). However, this overlooks developer choices made during data collection, data processing, and even the model training stages, all of which can influence the final model, as discussed above. Similarly, most existing studies define indistinguishability solely in terms of loss (Castillo et al., 2008; Teney et al., 2022; Paes et al., 2023; Du et al., 2024; Hamman et al., 2024), disregarding other crucial choices involved in designing evaluation criteria and model selection, as previously mentioned.
@@
(1) $\delta^{P}_{i}(h_{1},h_{2})\leq\epsilon^{P}_{i}\;\;\;\;\forall\;(\delta^{P}_{i},\epsilon^{P}_{i})\in(\mathbf{\Delta}^{P},\mathcal{E}^{P})$
We now turn to defining the context in which these models exhibit diverse behaviour, i.e., multiplicity. Again, as previously noted, most existing work focuses on either predictive or explanation multiplicity (see Figure 1), with limited attention given to other forms such as fairness multiplicity (Islam et al., 2021; Coston et al., 2021; Ganesh et al., 2023; Gillis et al., 2024; Sellam et al., [n. d.]; Wang and Russakovsky, 2021), OOD robustness multiplicity (Teney et al., 2022; McCoy et al., 2020), model complexity multiplicity (Semenova et al., 2024; Liu et al., 2022; Rudin et al., 2024; Semenova et al., 2022; Coupkova and Boutin, 2024), and feature interaction multiplicity (Li et al., [n. d.]), among others. To capture the various impacts discussed above, we generalize multiplicity by binding it to a metric delta $\delta^{M}$ and corresponding threshold $\epsilon^{M}$, on models belonging to the Rashomon set. Formally,
Definition 3.2 (Model Multiplicity).
@@
(2) $\delta^{P}_{i}(h_{1},h_{2})\leq\epsilon^{P}_{i}\;\;\;\;\forall\;(\delta^{P}_{i},\epsilon^{P}_{i})\in(\mathbf{\Delta}^{P},\mathcal{E}^{P})\quad\quad\textrm{and}\quad\delta^{M}(h_{1},h_{2})>\epsilon^{M}\end{aligned}$
Although our definitions are quite similar to those in existing literature, we have deliberately generalized them to encompass a broader range of underrepresented works. These are not radical changes, but we believe are crucial in drawing attention to various overlooked choices in the model design pipeline and to better understand their role in multiplicity. However, we also recognize that in certain contexts, a specific definition of multiplicity might be needed. In such contexts, our definition can be reduced appropriately to match the use case.
3.2. Evaluating Multiplicity
@@
Ambiguity (Marx et al., 2020) Predictive Multiplicity Multi-Class Classification Yes Dataset
Obscurity (Cavus and Biecek, 2024a) Predictive Multiplicity Multi-Class Classification No Dataset
Discrepancy (Marx et al., 2020) Predictive Multiplicity Multi-Class Classification Yes Dataset
Indistinguishability (Paes et al., 2023) Predictive Multiplicity Multi-Class Classification - Dataset
Degree of Underspecification (Teney et al., 2022) Predictive Multiplicity Multi-Class Classification Yes Dataset
Viable Prediction Range (Watson-Daniels et al., 2023b) Predictive Multiplicity Probabilistic Classification Yes Individual
Rashomon Capacity (Hsu and Calmon, 2022) Predictive Multiplicity Probabilistic Classification Yes Individual
Multi-target Ambiguity (Watson-Daniels et al., 2023a) Predictive Multiplicity Multi-Target Classification Yes Dataset
Rank List Sensitivity (Oh et al., 2022) Predictive Multiplicity Recommender Systems - Dataset
Std. of Scores (Heljakka et al., 2022; Long et al., 2024) Predictive Multiplicity Agnostic No Dataset
Representational Multiplicity (Heljakka et al., 2022) Procedural Multiplicity Agnostic No Dataset
Region Similarity Score (Somepalli et al., 2022) Procedural Multiplicity Agnostic - Dataset
Variance Arbitrariness Agnostic No Individual
Self-consistency (Cooper et al., 2024) Arbitrariness Multi-Class Classification No Individual
$\epsilon$-robust to Dataset Multiplicity (Meyer et al., 2023) Dataset Multiplicity Regression - Dataset
Unfairness Range (Mata et al., 2022) Fairness Multiplicity Agnostic Yes Dataset
Rashomon Ratio (Semenova et al., 2022) Size of Rashomon Set Agnostic Yes Dataset
Underspecification Score (Madras et al., 2019) Underspecification Multi-Class Classification No Individual
Accuracy Under Intervention (Ganesh, 2024) Metric Multiplicity Multi-Class Classification - Dataset
@@
Consistency (Watson et al., 2022) Explanation Multiplicity Agnostic No Dataset
Model Class Reliance (Fisher et al., 2019) Explanation Multiplicity Model Reliance (Fisher et al., 2019) Yes Dataset
Attribution Agreement (Krishna et al., 2022; Brunet et al., 2022) Explanation Multiplicity Feature Attribution - Individual
Profile Disparity Index (Kobylińska et al., 2024) Explanation Multiplicity Model Profile (Kobylińska et al., 2024) - Dataset
Inv. Cost of Neg. Surprise (Pawelczyk et al., 2020) Explanation Multiplicity Counterfactuals - Dataset
@@
Variable Importance Clouds (Dong and Rudin, 2019) Explanation Multiplicity Feature Attribution Yes Dataset
Coverage & Interval Width  (Marx et al., 2023) Explanation Uncertainty Agnostic No Individual
@@
Table 1. Existing multiplicity metrics in the literature, along with original objective, problem setting, monotonicity wrt to the Rashomon set, and metric resolution. Metrics with no entry under ‘Monotonic’ are defined to compare two models, and not the Rashomon set.
@@
To complement the expanded definitions, we compile a comprehensive list of metrics from the literature that measure various forms of multiplicity in Table 1. For each metric, we mark its original objective, the problem setting, resolution (i.e., whether the metric applies to individual data points or the entire dataset), and if it exhibits set monotonicity within the Rashomon set (i.e., whether a reduction in the Rashomon set size implies a monotonic change in the metric). Monotonicity in a metric is particularly desirable in certain scenarios (Hsu and Calmon, 2022), because it ensures that reducing the Rashomon set size—by improving specification—will either decrease or maintain the multiplicity. However, in other contexts, metrics that are not necessarily monotonic, such as variance, can be equally valuable. For instance, when it is important to quantify not just the range but also the dispersion of arbitrariness, metrics that capture such variability become crucial. Ultimately, the relevance of monotonicity in a metric depends on the specific context in which it is applied. By systematically recording different facets of each metric in Table 1, we hope to provide a structure for practitioners to identify the most suitable metric for their specific needs in evaluating multiplicity.
@@
Beyond the metrics themselves, we also revisit the challenge of training multiple models to evaluate multiplicity, exploring more efficient alternatives. Madras et al. (2019) propose a local ensembling technique to quantify underspecification by analyzing the loss curvature, eliminating the need to train multiple models. Hsu et al. (2024) use a similar local approach and show that Monte Carlo dropout can be adopted to approximate multiplicity when constrained by utility considerations (i.e., Rashomon sets). In contrast to these local methods, Kissel and Mentch (2024) introduce a model path selection technique that incrementally builds from simpler to more complex models. This approach efficiently constructs the Rashomon set by recursively increasing the complexity of plausible models. Other model class-specific techniques have also been proposed to explore Rashomon sets more efficiently (Hsu et al., [n. d.]; Mata et al., 2022; Zhong et al., 2024; Xin et al., 2022). Despite these advancements, much work remains to be done to make the enumeration of Rashomon sets more efficient and practical.
4. Multiplicity, Uncertainty and Bias-Variance Decomposition
@@
When examining arbitrariness in decision-making, machine learning research often focuses on prediction uncertainty (Gal, 2016; Gal et al., 2017; Kendall and Gal, 2017; Smith et al., [n. d.])–model’s lack of confidence in its predictions–or the bias-variance decomposition (Domingos, 2000; Kong and Dietterich, 1995; BREIMAN, 1996)–dividing the model error into how well the model fits the data (bias) and its sensitivity to changes in data (variance). With extensive literature already present in these areas, a natural question arises: What unique perspectives does multiplicity bring to the discussion of arbitrariness not already covered by these concepts? In this section, we formalize the interplay between multiplicity, uncertainty, and bias-variance decomposition, addressing this question mathematically and through practical recommendations for when different perspectives on arbitrariness are most valuable.
4.1. Multiplicity and Uncertainty
@@
Prediction Uncertainty: We start by defining uncertainty, drawing heavily from Gal (2016); Gal et al. (2017); Kendall and Gal (2017); Smith et al. ([n. d.]). Prediction uncertainty quantifies the degree of confidence—or lack thereof—in a model’s predictions. As it reflects the lack of confidence in a model’s predictions, uncertainty is often represented as the randomness (or entropy) in those predictions. Formally, prediction uncertainty is commonly defined as:
(3) $U(x,D)={\textrm{H}}_{y}[{\textrm{Prob}}(y|x,D)]$ $={\textrm{H}}_{y}[\Sigma_{\theta}[{\textrm{Prob}}(y|x,\theta)*{\textrm{Prob}}(\theta|D)]]$
(4) $={\textrm{H}}_{y}[{\textrm{E}}_{\theta\sim{\textrm{Prob}}(\theta|D)}[{\textrm{Prob}}(y|x,\theta)]]$
[…]
Prediction uncertainty is also typically divided into aleatoric and epistemic uncertainty (Hüllermeier and Waegeman, 2021; Senge et al., 2014; Kendall and Gal, 2017). Aleatoric uncertainty, connected to “distributional complexity” (Zhang et al., 2020), is the intrinsic difficulty of modelling the relationship in the data distribution, often described as the uncertainty left with access to infinite data. In contrast, epistemic uncertainty, connected to “approximation complexity” (Zhang et al., 2020), stems from a lack of knowledge about the best model to use, i.e., the challenge of accurately capturing the true data distribution. Epistemic uncertainty does not have a straightforward formulation. Instead, it is commonly defined as the uncertainty remaining after accounting for aleatoric uncertainty.
@@
For the interested reader, we encourage exploring the uncertainty literature (Hüllermeier and Waegeman, 2021; Tran et al., 2020; Kendall and Gal, 2017; Gal, 2016), as we do not expand on these concepts in our work. We simply restate these definitions to compare them directly with multiplicity.
@@
(7) $M(x,D)={\textrm{H}}_{y}[{\textrm{E}}_{\theta\sim{\textrm{Prob}}_{R}(\theta|D)}[{\textrm{Prob}}(y|x,\theta)]]\quad\quad\textrm{s.t.}\quad{\textrm{Prob}}_{R}(\theta|D)=\left\{\begin{array}[]{lr}0&\textrm{if}\;\theta\notin R\\ \dfrac{{\textrm{Prob}}(\theta|D)}{\sum_{\theta\in R}{\textrm{Prob}}(\theta|D)}&\textrm{if}\;\theta\in R\end{array}\right\}$
@@
where ${\textrm{Prob}}_{R}(\theta|D)$ represents a modified probability distribution that only includes the models in the Rashomon set $R$.
@@
Comparing equations 4 and 7, it is clear that the expectation terms are defined over different distributions: over all possible models for uncertainty (eq 4), and over only models within the Rashomon set for multiplicity (eq 7). But why does this distinction matter, and why should we care about both? To answer this, we discuss practical scenarios where viewing a problem through the lens of multiplicity is more appropriate than uncertainty and vice-versa.
@@
Uncertainty or Multiplicity? Choosing the Right Lens. Multiplicity examines prediction consistency, while uncertainty assesses confidence. Uncertainty is better suited for modelling the information-theoretic relationship between data and the predictions derived from it. Multiplicity, on the other hand, offers the most relevant perspective for actively exploring the various interpretations that can emerge during learning. Similarly, when examining how different modelling choices or model selection criteria can influence outcomes, the lens of multiplicity proves invaluable.
We outline some characteristics to look for when deciding whether to use the lens of multiplicity or uncertainty.
@@
  • Uncertainty provides an information-centric perspective. As uncertainty definitions are derived from information theory (Box and Tiao, 1973; Kendall and Gal, 2017; Gal, 2016), it is a fundamentally better fit for related analyses. For example, uncertainty plays a crucial role in active learning, by finding instances most likely to provide maximum new ‘information’ (Yang and Loog, 2016; Sharma and Bilgic, 2017; Nguyen et al., 2022).
@@
  • Uncertainty is sufficient when dealing with distributional complexity. Noise in real-world data may result in a lack of predictive power to make reliable decisions (Wang et al., 2024). Having access to different interpretations through multiplicity adds little value in such cases—multiple ways of delivering incorrect answers do not necessarily enhance usefulness.
@@
  • Uncertainty quantification can be more efficient, but research in multiplicity quantification is growing rapidly. Uncertainty is streamlined into modern machine learning pipelines through Bayesian networks and model calibration (Jospin et al., 2022; Niculescu-Mizil and Caruana, 2005; Vaicenavicius et al., 2019), offering a cost-effective alternative to multiplicity. Even when training multiple models, there is typically no definitive way to ensure that every trained model falls within the Rashomon set, and thus not all trained models contribute to measuring multiplicity whereas they are still valuable for quantifying uncertainty. That said, advancements in multiplicity research, as previously discussed in §3.2, have already been challenging this dynamic and may continue to reshape it further in the future (Madras et al., 2019; Hsu et al., 2024; Kissel and Mentch, 2024; Hsu et al., [n. d.]; Mata et al., 2022; Zhong et al., 2024).
@@
  • Multiplicity is aligned with learning theory and hierarchical optimization. Every decision in the learning algorithm influences the underlying optimization. Multiplicity can help scrutinize how each choice shapes the final model. Applications include the impact of data processing, random seeds, hyperparameters, etc. (Islam et al., 2021; Ganesh et al., 2023; Meyer et al., 2023; Ganesh, 2024; Cavus and Biecek, 2024b), or, broadly, any form of bi-level or constrained optimization, including meta-objectives (Black et al., 2024b; Semenova et al., 2024; Sun et al., [n. d.]; Ganesh et al., 2023; Gerchick et al., 2023).

### v2

@@ 2.2 Impact of Rashomon Effect
The most straightforward example of multiplicity is conflicting predictions from models in the Rashomon set, known as predictive multiplicity (Marx, Calmon, and Ustun 2020). Such conflicts create arbitrariness in decision-making, undermining the reliability of these models and hampering effective planning (D’Amour et al. 2022; Cooper, Frankle, and De Sa 2022; Watson-Daniels et al. 2024; Milani Fard et al. 2016). While predictive multiplicity can be harmful in critical domains, such as medical or legal decisions, it is, however, not inherently bad. For instance, purposefully controlled arbitrariness (called ‘randomness’ to distinguish from uncontrolled arbitrariness) can help address the concerns of outcome homogenization (Jain et al. 2024; Jain, Creel, and Wilson 2024; Creel and Hellman 2022; Barocas, Hardt, and Narayanan 2023).
@@
Unsurprisingly, predictive multiplicity has received significant attention in the literature ($35\%$ of papers in our systematic review; see Figure 1). Yet, this is only one aspect of the Rashomon effect. Consider, for instance, the inconsistency in explanations provided by models within the Rashomon set. Studies have shown that models in the Rashomon set often produce conflicting feature attribution scores (Laberge et al. 2023; Li, Barnard, and Deng 2024; Gunasekaran, Mistry, and Chen 2024; Müller et al. 2023; Poiret et al. 2023; Okazaki et al. 2024; Tan 2024; Kumar et al. 2023), which can undermine trust, e.g., confusing clinicians during AI-assisted diagnostics. Similarly, counterfactuals generated by one model in the Rashomon set often fail to transfer to others (Pawelczyk, Broelemann, and Kasneci 2020; Leofante, Botoeva, and Rajani 2023; Jiang et al. 2024a; Hasan and Talbert 2022). This poses significant challenges for algorithmic recourse when models are regularly updated, as recourse provided by one model may become invalid when replaced by another, questioning their legitimacy and undermining user trust (Rawal, Kamar, and Lakkaraju 2021; Leofante, Botoeva, and Rajani 2023).
@@
More broadly, multiplicity is any form of behavioural difference between models in the context of real-world consequences. These effects are multifaceted, and while predictive and explanation multiplicities are the most recognized, narrowing our focus underestimates the broader risks associated with other underrepresented forms of multiplicity (only $\sim 24\%$ of papers in our systematic review cover other forms of multiplicity; see Figure 1). In the next section, we will formally define both Rashomon sets and multiplicity, expanding upon existing definitions in the literature.
3 Definitions and Metrics
@@
We now define Rashomon sets and multiplicity, building on existing literature while expanding the scope to encompass a wider range of works. Additionally, we review multiplicity metrics and the literature on measuring multiplicity.
3.1 Formalizing Rashomon Sets and Multiplicity
@@
The concept of multiplicity is deeply rooted in the Rashomon effect, and the models illustrating this Rashomon effect are together known as a Rashomon set, a set of competing models, a set of good models, $\epsilon$-Rashomon set, $\epsilon$-Level set, etc. We’ll stick with the term Rashomon set for consistency. Rashomon set represents a set of models that are practically indistinguishable, underscoring the arbitrariness in choosing one model over another. Thus, we need to begin by defining these models, i.e., the Rashomon set.
@@
Existing work in multiplicity tends to adopt a narrow view of these models. Much of the research that formalizes the multiplicity problem restricts model choices to a specific hypothesis class $\mathcal{H}$ and/or limits them to training on a fixed, pre-processed dataset $\mathcal{D}$ (Marx, Calmon, and Ustun 2020; Teney, Peyrard, and Abbasnejad 2022; Watson-Daniels et al. 2023). However, this overlooks developer choices made during data collection, data processing, and even model training, all of which can influence the final model, as discussed above. Similarly, most existing studies define indistinguishability solely in terms of loss (Teney, Peyrard, and Abbasnejad 2022; Paes et al. 2023; Du, Ngo, and Wu 2024; Hamman et al. 2024), disregarding other choices involved in designing evaluation criteria and model selection.
@@
$\delta^{P}_{i}(h_{1},h_{2})\leq\epsilon^{P}_{i}\;\;\;\;\forall\;(\delta^{P}_{i},\epsilon^{P}_{i})\in(\mathbf{\Delta}^{P},\mathcal{E}^{P})$ (1)
Next, we turn to defining the context in which these models exhibit diverse behaviour, i.e., multiplicity. Again, as previously noted, most existing work focuses on either predictive or explanation multiplicity (see Figure 1), with limited attention given to other forms such as fairness multiplicity (Islam, Pan, and Foulds 2021; Coston, Rambachan, and Chouldechova 2021; Ganesh et al. 2023; Gillis, Meursault, and Ustun 2024; Wang and Russakovsky 2021), allocation multiplicity (Jain et al. 2025), OOD robustness multiplicity (Teney, Peyrard, and Abbasnejad 2022; McCoy, Min, and Linzen 2020), model complexity multiplicity (Semenova et al. 2024; Liu et al. 2022; Rudin et al. 2024; Semenova, Rudin, and Parr 2022; Coupkova and Boutin 2024), and feature interaction multiplicity (Li et al. 2024), among others. To capture the various impacts, we generalize multiplicity by binding it to a metric delta $\delta^{M}$ and corresponding threshold $\epsilon^{M}$, on models belonging to the Rashomon set. Formally,
Definition 2 (Model Multiplicity)
@@
$\begin{aligned} &\delta^{P}_{i}(h_{1},h_{2})\leq\epsilon^{P}_{i}\;\;\;\;\forall\;(\delta^{P}_{i},\epsilon^{P}_{i})\in(\mathbf{\Delta}^{P},\mathcal{E}^{P})\\ \textrm{and}\quad&\delta^{M}(h_{1},h_{2})>\epsilon^{M}\end{aligned}$ (2)
Although our definitions are quite similar to existing literature, we have deliberately generalized them to encompass a broader range of underrepresented works. These are not radical changes, but we believe are crucial in drawing attention to various overlooked choices in model design and to better understand their role in multiplicity. However, we also recognize that in certain contexts, a specific definition of multiplicity might be needed. In such contexts, our definition can be reduced appropriately to match the use case.
3.2 Evaluating Multiplicity
@@
Ambiguity Predictive Multiplicity Multi-Class Classification Yes Dataset
Obscurity Predictive Multiplicity Multi-Class Classification No Dataset
Discrepancy Predictive Multiplicity Multi-Class Classification Yes Dataset
Degree of Underspecification Predictive Multiplicity Multi-Class Classification Yes Dataset
Viable Prediction Range Predictive Multiplicity Probabilistic Classification Yes Individual
Rashomon Capacity Predictive Multiplicity Probabilistic Classification Yes Individual
Multi-target Ambiguity Predictive Multiplicity Multi-Target Classification Yes Dataset
Rank List Sensitivity Predictive Multiplicity Recommender Systems - Dataset
Std. of Scores Predictive Multiplicity Agnostic No Dataset
Self-consistency Arbitrariness Multi-Class Classification No Individual
Representational Multiplicity Procedural Multiplicity Agnostic No Dataset
Region Similarity Score Procedural Multiplicity Agnostic - Dataset
$\epsilon$-robust to Dataset Multiplicity Dataset Multiplicity Regression - Dataset
Unfairness Range Fairness Multiplicity Agnostic Yes Dataset
Rashomon Ratio Size of Rashomon Set Agnostic Yes Dataset
Underspecification Score Underspecification Multi-Class Classification No Individual
Accuracy Under Intervention Metric Multiplicity Multi-Class Classification - Dataset
@@
Consistency Explanation Multiplicity Agnostic No Dataset
Model Class Reliance Explanation Multiplicity Model Reliance Yes Dataset
Attribution Agreement Explanation Multiplicity Feature Attribution - Individual
Profile Disparity Index Explanation Multiplicity Model Profile - Dataset
Inv. Cost of Neg. Surprise Explanation Multiplicity Counterfactuals - Dataset
@@
Variable Importance Clouds Explanation Multiplicity Feature Attribution Yes Dataset
Coverage & Interval Width Explanation Uncertainty Agnostic No Individual
@@
Table 1: Multiplicity metrics along with original objective, problem setting, monotonicity wrt to the Rashomon set, and metric resolution. Metrics with no entry under ‘Monotonic’ compare only two models. Metric citations in the Appendix (§B).
@@
We compile a comprehensive list of metrics from the literature that measure various forms of multiplicity in Table 1. For each metric, we mark its original objective, the problem setting, resolution (i.e., whether the metric applies to individual data points or the entire dataset), and if it exhibits set monotonicity within the Rashomon set (i.e., whether a reduction in the Rashomon set size implies a monotonic change in the metric). Monotonicity in a metric can be desirable in certain scenarios (Hsu and Calmon 2022), because it ensures that reducing the Rashomon set size will either decrease or maintain the multiplicity. By systematically recording different facets of each metric in Table 1, we hope to provide a structure for practitioners to identify the most suitable metric for their specific needs.
@@
We also revisit the challenge of training multiple models to evaluate multiplicity, exploring more efficient alternatives. Madras, Atwood, and D’Amour (2019) propose a local ensembling technique to quantify underspecification by analyzing the loss curvature, eliminating the need to train multiple models. Hsu et al. (2024b) use a similar local approach and show that Monte Carlo dropout can be adopted to approximate multiplicity when constrained by utility considerations (i.e., Rashomon sets). In contrast to these local methods, Kissel and Mentch (2024) introduce a model path selection technique that incrementally builds from simpler to more complex models. This approach efficiently constructs the Rashomon set by recursively increasing the complexity of plausible models. Other model class-specific techniques have also been proposed to explore Rashomon sets more efficiently (Hsu et al. 2024a; Mata, Kanamori, and Arimura 2022; Zhong et al. 2024; Xin et al. 2022). Despite these advancements, much work remains to be done to make the enumeration of Rashomon sets more efficient and practical.
4 Multiplicity, Uncertainty and Bias-Variance Decomposition
@@
When examining arbitrariness in decision-making, machine learning research often focuses on prediction uncertainty (Gal 2016; Smith et al. 2024) or the bias-variance decomposition (Domingos 2000; Kong and Dietterich 1995; Brieman 1996). With extensive literature already present in these areas, a natural question arises: What unique perspectives does multiplicity bring that is not already covered by these concepts? In this section, we formalize the interplay between multiplicity, uncertainty, and bias-variance decomposition, addressing this question both mathematically and through practical recommendations.
4.1 Multiplicity and Uncertainty
@@
Prediction Uncertainty: We start by defining uncertainty, drawing heavily from Smith et al. (2024). Prediction uncertainty quantifies the degree of confidence–or lack thereof–in a model’s predictions. As it reflects the lack of confidence in a model’s predictions, uncertainty is often represented as the randomness (or entropy) in those predictions. Formally, prediction uncertainty is commonly defined as:
$U(x,D)$ $={\textrm{H}}_{y}[{\textrm{Prob}}(y|x,D)]$ (3)
$={\textrm{H}}_{y}[{\textrm{E}}_{\theta\sim{\textrm{Prob}}(\theta|D)}[{\textrm{Prob}}(y|x,\theta)]]$ (4)
where $x$ is the input for which we’re measuring uncertainty, $y$ is the output, $D$ is the training data, and $\theta$ is the parametric representation of models. $U(x,D)$ is the prediction uncertainty, while ${\textrm{H}}[.],{\textrm{E}}[.],{\textrm{Prob}}[.]$ represents entropy, expectation, and probability distribution, respectively. The subscript for each statistical measure specifies the random variable or the distribution on which the measure is calculated.
@@
We redirect the interested reader to the uncertainty literature (Hüllermeier and Waegeman 2021; Tran, Snoek, and Lakshminarayanan 2020; Kendall and Gal 2017; Gal 2016), as we do not expand further here. We simply restate these definitions to compare them with multiplicity.
@@
$M(x,D)={\textrm{H}}_{y}[{\textrm{E}}_{\theta\sim{\textrm{Prob}}_{R}(\theta)}[{\textrm{Prob}}(y|x,\theta)]]$ (5)
${\textrm{Prob}}_{R}(\theta)=\dfrac{{\textrm{Prob}}(\theta|D)}{\sum\limits_{\theta\in R}{\textrm{Prob}}(\theta|D)}\;\;\textrm{if}\;\theta\in R;\;\;0\;\textrm{otherwise}$ (6)
@@
where ${\textrm{Prob}}_{R}(\theta)$ is a modified probability distribution that only includes the models in the Rashomon set $R$.
@@
Comparing equations 4 and 5, it is clear that the expectation terms are defined over different distributions: over all possible models for uncertainty (eq 4), and over only models within the Rashomon set for multiplicity (eq 5). In other words, while uncertainty measures the potential variance in model predictions for the complete hypothesis class, multiplicity focuses only on models in the Rashomon set, i.e., has a more practical view towards model selection. But why does this distinction matter, and why should we care about both? To answer this, we discuss practical scenarios where viewing a problem through the lens of multiplicity is more appropriate than uncertainty and vice versa.
@@
Uncertainty or Multiplicity? Choosing the Right Lens. Multiplicity examines prediction consistency, while uncertainty assesses confidence. Uncertainty is better suited for modelling the information-theoretic relationships, while multiplicity, on the other hand, actively explores the various interpretations that can emerge during learning. Similarly, when examining how different modelling choices or model selection criteria can influence outcomes, the lens of multiplicity proves invaluable. We outline some characteristics to look for when deciding between the two.
@@
  • Uncertainty provides an information-centric perspective. As uncertainty definitions are derived from information theory (Box and Tiao 1973; Kendall and Gal 2017; Gal 2016), it is a fundamentally better fit for related analyses. For example, uncertainty plays a crucial role in active learning, by finding instances most likely to provide maximum new ‘information’ (Yang and Loog 2016; Sharma and Bilgic 2017; Nguyen, Shaker, and Hüllermeier 2022).
@@
  • Uncertainty is sufficient for distributional complexity. Noise in real-world data may result in a lack of predictive power to make reliable decisions (Wang et al. 2024). Having access to different interpretations through multiplicity adds little value in such cases.
@@
  • Uncertainty quantification is more efficient, but research in multiplicity quantification is growing rapidly. Uncertainty is streamlined in modern machine learning pipelines through Bayesian networks and model calibration (Niculescu-Mizil and Caruana 2005; Vaicenavicius et al. 2019), offering a cost-effective alternative to multiplicity. Even when training multiple models, there is no definitive way to ensure that every trained model falls in the Rashomon set. Thus, not all trained models contribute to measuring multiplicity, whereas they are still valuable for quantifying uncertainty. That said, advancements in multiplicity research are already challenging this dynamic (§3.2) and may continue to reshape it in the future.
@@
  • Multiplicity is aligned with learning theory and hierarchical optimization. Every decision in the learning algorithm influences the underlying optimization. Multiplicity can help scrutinize how each choice shapes the final model. Applications include the impact of data processing, random seeds, hyperparameters, etc. (Islam, Pan, and Foulds 2021; Meyer, Albarghouthi, and D’Antoni 2023; Ganesh 2024; Cavus and Biecek 2024b), or, broadly, any form of bi-level or constrained optimization (Black et al. 2024; Semenova et al. 2024; Sun, Wang, and Rudin 2024; Ganesh et al. 2023; Gerchick et al. 2023).